Find the equation for the circle with a diameter whose endpoints are (4,-1) and (2,4)
Answer by lwsshak3(11628)

Find the equation for the circle with a diameter whose endpoints are (4,-1) and (2,4)
**
standard form of equation for a circle: (x-h)^2+(y-k^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
center=midpoint of diameter
=(2+4)/2,(4-1)/2
=(3,3/2)
equation: (x-3)^2+(y-3/2)^2=r^2
Solve for r^2 using coordinates of one of given points(2,4) on the circle.
(2-3)^2+(4-3/2)^2=r^2
1+25/4=r^2
r^2=29/4
equation: (x-3)^2+(y-3/2)^2=29/4